# Geometric Solution of the Loop Equation and Quark Confinement in Planar QCD

### v2

@@ 11 Topological String with Nonlinear Endpoint Dynamics
This system describes a topological field theory: the action reduces entirely to boundary values as a consequence of the imposed Cauchy-Riemann equations. Physically, this action represents the net area of two conformal maps, $f_{1}(\mathbb{D})$ and $f_{2}(\mathbb{D})$, with no remaining bulk degrees of freedom. In string theory language, there are no vibrations of the string body; there are only the nonlinear dynamics of its endpoints determined by the loop shape. In this sense, it is analogous to 2D quantum gravity, where the Einstein-Hilbert action reduces to the Euler characteristic, independent of the local geometry in the bulk.
@@
This corresponds to setting an optimal isometry of the boundary condition rather than averaging over these isometries. In terms of the partition function $Z=\exp\left(-\kappa S[RC]\right)$, our prescription selects the minimal value, rather than performing a saddle point integration around the maximal value of this factor.
Let us stress that we are solving the classical loop equations of large-$N_{c}$ QCD rather than quantizing a string theory. The internal degrees of freedom found in non-critical string theory (such as Polyakov’s Liouville mode) do not apply here, as our theory is based on the classical dynamics of Wilson loop trace functionals in planar QCD. We utilize string theory as a mathematical tool, rather than as a physical object describing the condensation of electric flux in the QCD vacuum.
In our picture, the fluctuations of this electric flux are described by planar graphs for gluons in another factor $W_{pert}[C]$. The confining factor $Z[C]$ only reflects the mean properties of the vacuum, frozen in the large $N_{c}$ limit. The optimal rotation $R_{*}$ described in section 5.1, aligning the SD/ASD projector with the highest plane of the Gram basis determines the boundary conditions for the holomorphic string:
@@ 12 String Scale $\kappa$ and Gluon Condensate
The string mass scale $\kappa$ is not fixed by the MM equation itself. As we have seen, the confining factor $\exp\left(-\kappa S[C]\right)$ passes through the equation without correlation with the perturbative solution $W_{pert}[C]$. This factor reflects the non-perturbative properties of the QCD vacuum rather than the perturbative fluctuations of the gluon field.
The fundamental non-perturbative property of the QCD vacuum is the gluon condensate [Shifman:1978bx]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$. According to our solution, $S[C]=2\sqrt{2}\pi r^{2}$. We compute the second area derivative at two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$ on the circle.
@@
The divergence scaling as $1/r^{2}$ arises from the second area derivative as a direct consequence of the scaling property of our Hodge minimal area. This is a nonlinear, nonlocal, homogeneous functional of degree two. Every area derivative involves two functional derivatives with respect to $\dot{C}$, lowering the degree by two. Thus, the second area derivative has degree minus two, inevitably scaling as $1/r^{2}$. The logarithmic divergence comes from the high-frequency spectral sum.
@@
where the constant $a\sim 1$ is calculable from our solution, and renormalization constant $Z$ for the area derivative operator applied to $W_{pert}[C]$ is calculable by the methods of [KorchemskyRadyushkin1987]. Presumably, the renormalization factor $Z^{2}$ on the LHS compensates the UV divergence $(\log n_{max})^{2}$ on the RHS, rendering the relation finite and renormalization-group invariant.
The function $f(z)=ze^{-z}$ (where $z=2\sqrt{2}\pi r^{2}\kappa$) reaches its maximum at $z=1$. This implies that no solutions exist outside this parameter region, suggesting a critical point. We identify this maximum as the natural matching point:
@@
Our association of $\kappa$ with the gluon condensate supports the physical picture of condensed electric flux originally conjectured by Nambu [Nambu:1974zg]. One may view the term $\kappa S[C]$ as the action cost of a field configuration with energy density $\langle G^{2}\rangle\sim\Lambda_{QCD}^{4}$ spread over a “pancake” surface of thickness $\Lambda_{QCD}^{-2}$. This estimate interprets the dressing factor $\exp\left(-\kappa S[C]\right)$ as a statistical probability $\exp\left(-\beta\delta E\right)$ for a large quark loop propagating through the QCD vacuum (the dual Meissner effect).
@@ 13 Conclusion
We have presented an exact solution for the quaternionic Hodge–dual minimal surface, a new class of minimal surfaces in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ that is exactly solvable by reduction to free conformal maps. The minimized area is the non-trivial spectral invariant $S_{\min}(C)=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})$, where $\{\lambda_{i}\}$ are the eigenvalues of the Douglas–Gram matrix $G$. This result is mathematically robust: the optimal surface is determined by the two dominant Douglas eigenvectors, and the underlying worldsheet theory is quantum-exact.
This Hodge–dual minimal area $S[C]$ enters the non-perturbative factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop. This factor is an exact zero mode factor satisfying the MM equation in QCD [MMEq79, MM1981NPB, Mig83]. The QCD OPE provides the estimate (12.8) relating our mass scale $\kappa$ to the gluon condensate and QCD mass scale. The primary physical implication is a first-principles geometric proof of confinement, as stated in our Confinement Theorem (Theorem 9.1). This theorem provides an explicit, shape-dependent formula for the effective string tension $\sigma_{\mathrm{eff}}[C]$ (defined as a ratio of $S[C]$ to the Euclidean area $A[C]$). This parameter depends on the loop shape and is bounded from below as $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$. This solution lays the groundwork for a quantitative theory of the high-level hadron spectrum.

### v3

@@ 10 Topological String with Nonlinear Endpoint Dynamics
This system describes a topological field theory: the action reduces entirely to boundary terms due to the imposed Cauchy-Riemann equations. Physically, this action represents the net area of two conformal maps, $f_{1}(\mathbb{D})$ and $f_{2}(\mathbb{D})$, with no remaining bulk degrees of freedom. In string theory language, there are no vibrations of the string body; there are only the nonlinear dynamics of its endpoints determined by the loop shape. In this sense, it is analogous to 2D quantum gravity, where the Einstein-Hilbert action reduces to the Euler characteristic, independent of the bulk geometry.
@@
This prescription corresponds to setting an optimal isometry of the boundary condition rather than averaging over these isometries. In terms of the partition function $Z=\exp\left(-\kappa S[RC]\right)$, our prescription selects the minimal value, rather than performing a saddle point integration around the maximal value of this factor.
Let us stress that we are solving the QCD loop equations rather than quantizing a fundamental string theory. The internal degrees of freedom found in non-critical string theory (such as Polyakov’s Liouville mode) do not apply here. We utilize string theory as a mathematical tool, acting as a computational dual. This approach mirrors the logic of the Eguchi-Kawai reduction [8], where auxiliary variables (quenched momenta) are introduced not as physical degrees of freedom, but as instruments to satisfy the loop equations.
In our picture, the fluctuations of this electric flux are described by gluon graphs in another factor $W_{pert}[C]$. The confining factor $Z[C]$ only reflects the mean properties of the vacuum.
The optimal rotation $R_{\ast}$ described in section 5.1, aligning the SD/ASD projector with the highest plane of the Gram basis, determines the boundary conditions for the holomorphic string:
@@ 11 String Scale $\kappa$ and Gluon Condensate
The string mass scale $\kappa$ is not fixed by the MM equation itself. As we have seen, the confining factor $\exp\left(-\kappa S[C]\right)$ passes through the equation without correlation with the perturbative solution $W_{pert}[C]$. This factor reflects the nonperturbative properties of the QCD vacuum rather than the perturbative fluctuations of the gluon field.
The fundamental nonperturbative property of the QCD vacuum is the gluon condensate [Shifman:1978bx]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$. According to our solution, $S[C]=2\sqrt{2}\pi r^{2}$. We compute the second area derivative at two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$ on the circle.
@@
The divergence scaling as $1/r^{2}$ arises from the second area derivative as a direct consequence of the scaling property of our Hodge minimal area. This area is a nonlinear, nonlocal, homogeneous functional of degree two. Every area derivative involves two functional derivatives with respect to $\dot{C}$, lowering the degree by two. Thus, the second area derivative has degree minus two, inevitably scaling as $1/r^{2}$. The logarithmic divergence comes from the high-frequency spectral sum.
@@
where the constant $a\sim 1$ is calculable from our solution, and renormalization constant $Z$ for the area derivative operator applied to $W_{pert}[C]$ is calculable by the methods of [KorchemskyRadyushkin1987]. Presumably, the renormalization factor $Z^{2}$ on the LHS compensates for the UV divergence $(\log n_{max})^{2}$ on the RHS, rendering the relation finite and renormalization-group invariant. The Fourier cutoff $n_{max}$ can be related to the UV cutoff $\Lambda$ of QCD by estimating the sides of the polygonal loop with discrete Fourier transform instead of continuum one $\Delta\theta=2\pi/n_{max}\sim 1/(\Lambda|C^{\prime}|)\sim 1/(\Lambda r)$, so that $n_{max}\sim\Lambda r$.
The function $f(z)=ze^{-z}$ (where $z=2\sqrt{2}\pi r^{2}\kappa$) reaches maximum at $z=1$. This limitation implies that no solutions exist outside this parameter region, suggesting the presence of a critical point. We identify this maximum as the natural matching point:
@@
Our association of $\kappa$ with the gluon condensate supports the physical picture of condensed electric flux originally conjectured by Nambu [Nambu:1974zg]. One may view the term $\kappa S[C]$ as the action cost of a field configuration with energy density $\langle G^{2}\rangle\sim\Lambda_{QCD}^{4}$ spread over a “pancake” surface of thickness $\Lambda_{QCD}^{-2}$ containing the 44- volume $\Lambda_{QCD}^{-2}S[C]$. This estimate interprets the dressing factor $\exp\left(-\kappa S[C]\right)$ as a statistical probability $\exp\left(-\beta\delta E\right)$ for a large quark loop propagating through the QCD vacuum (the dual Meissner effect).
@@ 12 Conclusion
We have presented an exact solution for the quaternionic Hodge–dual minimal surface, a new class of minimal surfaces in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ that can be solved exactly by reduction to free conformal maps. The minimized Hodge–dual area is the nontrivial spectral invariant:
$S[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$
where $\{\lambda_{i}\}$ are the eigenvalues of the Douglas–Gram matrix $G$. The two dominant Douglas eigenvectors determine the optimal surface, and the bulk dynamics reduces to a nonlinear boundary functional, depending on the loop geometry. There are no physical degrees of freedom associated with fluctuations of the shape of the surface, unlike ordinary string theory.
The Hodge–dual minimal area $S[C]$ enters the nonperturbative dressing factor $\exp\left(-\kappa S[C]\right)$ multiplying the perturbative Wilson loop. Thanks to the self–dual/anti–self–dual area derivative and the strict additivity of $S[C]$ at intersections, this dressing defines an exact zero mode of the loop–space diffusion operator. By Lemma 8.1, the multiplicative dressing preserves the entire MM multiloop hierarchy for any finite number of colors $N_{c}$, including the physical case $N_{c}=3$. Thus, the same geometric confining area law applies both in the planar limit and at finite $N_{c}$. The QCD OPE analysis provides the estimate (11.8), relating the mass scale $\kappa$ to the gluon condensate and the QCD scale.
The central physical implication is summarized in our Confinement Theorem (Theorem 9.1), which yields an explicit, shape–dependent formula for the effective string tension:
$\sigma_{\mathrm{eff}}[C]:=\frac{S[C]}{A_{\mathrm{geom}}[C]},$
defined as the ratio of the Hodge–dual area $S[C]$ to the Euclidean geometric area $A_{\mathrm{geom}}[C]=\mathrm{tr}\,G$. The effective tension depends on the loop geometry but obeys the universal lower bound $\sigma_{\mathrm{eff}}[C]\geq\sqrt{2}$, providing a rigorous geometric confinement criterion. Together with the OPE matching, this framework yields a confining area–law solution of the loop equations in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory and lays the basis for a computation of the high–lying meson spectrum.